# Unraveling overoptimism and publication bias in ML-driven science

### v1

@@ Abstract
Machine Learning (ML) is increasingly used across many disciplines with impressive reported results across many domain areas. However, recent studies suggest that the published performance of ML models are often overoptimistic and not reflective of true accuracy were these models to be deployed. Validity concerns are underscored by findings of a concerning inverse relationship between sample size and reported accuracy in published ML models across several domains. This is in contrast with the theory of learning curves in ML, where we expect accuracy to improve or stay the same with increasing sample size. This paper investigates the factors contributing to overoptimistic accuracy reports in ML-based science, focusing on data leakage and publication bias. Our study introduces a novel stochastic model for observed accuracy, integrating parametric learning curves and the above biases. We then construct an estimator based on this model that corrects for these biases in observed data. Theoretical and empirical results demonstrate that this framework can estimate the underlying learning curve that gives rise to the observed overoptimistic results, thereby providing more realistic performance assessments of ML performance from a collection of published results. We apply the model to various meta-analyses in the digital health literature, including neuroimaging-based and speech-based classifications of several neurological conditions. Our results indicate prevalent overoptimism across these fields and we estimate the inherent limits of ML-based prediction in each domain.
@@ Appendix C Extended results
Figure 5: Extended results based on Experiment 1 and for 5 additional cases introduced in Page 2.

### v2

@@ Abstract
Machine Learning (ML) is increasingly used across many disciplines with impressive reported results across many domain areas. However, recent studies suggest that the published performance of ML models are often overoptimistic. Validity concerns are underscored by findings of an inverse relationship between sample size and reported accuracy in published ML models, contrasting with the theory of learning curves where accuracy should improve or remain stable with increasing sample size. This paper investigates factors contributing to overoptimistic accuracy reports in ML-driven science, focusing on data leakage and publication bias. We introduce a novel stochastic model for observed accuracy, integrating parametric learning curves and the aforementioned biases. We then construct an estimator that corrects for these biases in observed data. Theoretical and empirical results show that our framework can estimate the underlying learning curve, providing realistic performance assessments from published results. Applying the model to meta-analyses in ML-driven science, including neuroimaging-based and speech-based classifications of neurological conditions, we find prevalent overoptimism and estimate the inherent limits of ML-based prediction in each domain.
@@ Appendix C Extended results
Figure 5: Extended results based on Experiment 1 and for 5 additional cases introduced in Section C.